# Towards Verifying Exact Conditions of Density Functional Theory Approximations

### v1

@@ VI-A Improving scalability of the solver
XCVerifier was unable to verify any of the exact conditions of the SCAN functional, which has been designed to satisfy all known exact conditions. SCAN is significantly more complex than the other functionals we considered, and also involves the use of transcendental functions such as $\exp$ and $\log$. This causes dReal to timeout even when the input domain is reduced. It would be interesting to investigate approaches to improve the performance of the solver so that it can tackle the SCAN functional. Apart from its popularity, the SCAN functional will serve as a fascinating use case: there are a progression of DFAs—rSCAN, r++SCAN, r2SCAN, r4SCAN—proposed in the literature that were designed with different adherence to exact conditions to improve the numerical stability of the original SCAN functional [73, 74, 75].
VI-B Expanding to more DFA/condition pairs
Our evaluation demonstrated the robustness of the XCEncoder to handle a wide variety of DFAs and exact conditions. The ultimate goal of our research is to be able to analyze all the 500+ functionals in Libxc for all known DFT exact conditions. Future work will continue to expand our evaluation, and will aim to integrate our verification tool into the Libxc, e.g., as part of the continuous integration (CI) for Libxc.
@@ VI-C Numerical issues with DFAs
Apart from verifying known exact conditions for DFA implementations, it would interesting to analyze numerical issues of the implementations with the goal of using formal methods to find and fix numerical issues in DFA implementations. This is a challenging problem involving reasoning about floating points and dealing with transcendental functions like $\sin$, $\log$ and $\exp$. The functional forms of DFAs themselves can also be a source of numerical issues. Some DFAs include different functions that apply to different input domains, and must ensure continuity when switching from one domain to another. Additionally, the parametrization of the DFA may cause issues. Even in the simple case of the Local Density Approximation (LDA), the Perdew-Zunger [76] parametrization of the results of Ceperley and Alder [77] includes potentially inaccurate numerical constants that lead to discontinuities of the exchange-correlation energy at a given matching point.
The functional form of a DFA may also make it sensitive to inaccuracies in its input data. While a given implementation of a DFA may yield correct answers for an exactly known (e.g. exponential) density, it may result in large numerical errors if the input density is noisy or if the density and its gradient are not numerically consistent. This is particularly problematic in regions of low density, e.g. far from a molecule placed in vacuum. Such large errors may lead to inaccurate energies or slow convergence in the solution of the Kohn-Sham equations. For example, the sensitivity of the SCAN functional requires the use of extremely fine grids to represent the electron density in order to avoid large numerical errors. This led some authors to modify the SCAN functional to avoid this numerical issue, resulting in a slightly different DFA [74]. In other cases, an analytical reformulation of a DFA is used to avoid numerical issues [78] without modifying it. These fixes are however ad hoc, and there is no known general recipe for avoiding the numerical issues of a DFA.
@@ VII Conclusion
This paper presented XCVerifier, an approach for verifying whether a DFA implementation satisfies DFT exact conditions. XCVerifier automatically encodes the DFA implementation from Libxc and a given exact condition into a dReal formula, symbolically performing any required derivative and limit calculations. It uses the dReal solver to verify whether the condition is true or find a violation to it for a given input domain. XCVerifier also implements a domain-splitting technique to improve performance by reducing solver timeouts, and isolate the input regions where the condition is satisfied or violated.
We evaluated XCVerifier by verifying seven exact conditions (from Pederson and Burke [40]) for five popular DFAs. XCVerifier was successfully able to verify or find a counterexample for 11 out of the 29 (valid) DFA/condition pairs, and it was able to verify all but a small portion of the input domain for an additional 7 pairs. However, it timed out for 11 pairs, which included all the conditions for the SCAN functional. We also found that the results of the PB approach, which used grid-search to check DFT exact conditions, were inconsistent with those of XCVerifier for 5 DFA/condition pairs. These results demonstrate promise and future challenges of using formal methods for DFT.

### v5

@@ VI-A Improving Scalability of the Solver
In our experiments, XCVerifier was unable to verify any of the exact conditions of the SCAN functional, which has been designed to satisfy all known exact conditions. SCAN is significantly more complex than the other functionals we considered, and also involves the use of transcendental functions such as $\exp$ and $\log$. This causes dReal to time out even for the relatively simple $E_{\rm c}$ non-positivity condition, and even when the input domain is reduced 32×32$\times$32 ×. It would be interesting to investigate approaches to improve the performance of the solver so that it can tackle the SCAN functional. Apart from its popularity, the SCAN functional will serve as a fascinating use case: there is a progression of DFAs—rSCAN, r++SCAN, r2SCAN, r4SCAN—proposed in the literature that were designed with different adherence to exact conditions to improve the numerical stability of the original SCAN functional [72, 73, 74].
VI-B Expanding to More DFA-Condition Pairs
Our evaluation demonstrated the robustness of the XCEncoder to handle a wide variety of DFAs and exact conditions. The ultimate goal of our research is to be able to analyze all the 500+ functionals in Libxc for all known DFT exact conditions. Future work will continue to expand our evaluation, and will aim to integrate our verification tool into Libxc, e.g., as part of the continuous integration (CI) for Libxc.
@@ VI-C Numerical Issues With DFAs
Apart from verifying known exact conditions for DFA implementations, it would be interesting to analyze numerical issues of the implementations with the goal of using formal methods to find and fix numerical issues in DFA implementations. This is a challenging problem involving reasoning about floating points and dealing with transcendental functions like $\sin$, $\log$ and $\exp$. The functional forms of DFAs themselves can also be a source of numerical issues. Some DFAs include different functions that apply to different input domains, and must ensure continuity when switching from one domain to another. Additionally, the parametrization of the DFA may cause issues. Even in the simple case of the Local Density Approximation (LDA), the Perdew-Zunger [75] parametrization of the results of Ceperley and Alder [76] includes potentially inaccurate numerical constants that lead to discontinuities of the exchange-correlation energy at a given matching point.
The functional form of a DFA may also make it sensitive to inaccuracies in its input data. While a given implementation of a DFA may yield correct answers for an exactly known (e.g., exponential) density, it may result in large numerical errors if the input density is noisy, or if the density and its gradient are not numerically consistent. This is particularly problematic in regions of low density, e.g., a point far from a molecule placed in vacuum. Such large errors may lead to inaccurate energies or slow convergence in the solution of the Kohn-Sham equations. For example, the sensitivity of the SCAN functional requires the use of extremely fine grids to represent the electron density in order to avoid large numerical errors. This led some authors to modify the SCAN functional to avoid this numerical issue, resulting in a slightly different DFA [73]. In other cases, an analytical reformulation of a DFA is used to avoid numerical issues [77] without modifying it. However, these fixes are ad hoc, and there is no known general recipe for avoiding the numerical issues of a DFA.
@@ VII Conclusion
This paper presented XCVerifier, an approach for verifying whether a DFA implementation satisfies the DFT exact conditions. XCVerifier automatically encodes the DFA implementation from Libxc and a given exact condition into a dReal formula, symbolically performing any required derivative calculations. It uses the dReal solver to verify whether the condition is true or find a violation to it for a given input domain. XCVerifier also implements a domain-splitting technique to improve performance, reducing solver timeouts, and isolating the input regions where the condition is satisfied or violated.
We evaluated XCVerifier by verifying seven exact conditions (from Pederson and Burke [28]) for five popular DFAs. XCVerifier was successfully able to verify or find a counterexample for 13 out of the 31 (valid) DFA-condition pairs, and it was able to partially verify an additional seven pairs. However, it timed out for 11 pairs, which included all the conditions for the SCAN functional. We found that the results of the PB approach, which used grid search to check DFT exact conditions, were largely consistent with those of XCVerifier. These results demonstrate promise and future challenges of using formal methods for DFT.